OpenStudy (anonymous):

first simplify equation then expression 12 over 20 and then 3x+12 over 5x+20

OpenStudy (saifoo.khan):

what's in between?

OpenStudy (anonymous):

12 over 20 is equation

OpenStudy (anonymous):

then 3x=12 over 5x+20 is the expression

OpenStudy (anonymous):

and I am sorry....12 over 20 is fraction

OpenStudy (anonymous):

\[\frac{12}{20}\times \frac{3x+12}{5x+20}\]

OpenStudy (anonymous):

i bet anyway

OpenStudy (saifoo.khan):

betting is not allowed in islam! :D

OpenStudy (anonymous):

haram this mister...

OpenStudy (anonymous):

i will have a drink instead!

OpenStudy (saifoo.khan):

lol, right.

OpenStudy (anonymous):

@jadams is this two different problems or one problem?

OpenStudy (anonymous):

\[\frac{12}{20}=\frac{4\times 3}{4\times 5}=\frac{3}{5}\]

OpenStudy (anonymous):

two separate problems but counted as one

OpenStudy (anonymous):

\[\frac{3x+12}{5x+20}=\frac{3(x+4)}{5(x+4)}=\frac{3}{5}\]